# Training on test proteins improves fitness, structure, and function prediction

### v1

@@ A.3.3 Models
Light attention + TTT.
When applying TTT to Light attention, we treat the frozen ESM-1b as the backbone $f$, the language modeling head of ESM-1b as the self-supervised head $g$, and the Light attention block as the fine-tuning head $h$.
@@
[] Learning rate Batch size Grad. acc. steps TTT steps LoRA rank $r$ LoRa $\alpha$ Fitness prediction ESM2 (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32, 64 30 - - ESM2 (650M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (650M) + TTT 4e-5, 4e-4, 4e-3 2, 4 4, 8, 16, 32 30 - - Structure prediction ESMFold + TTT 4e-4 4 4, 8, 32, 64 30 (max pLDDT) 4, 8, 32 8, 16, 32 ESM3 + TTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (max pLDDT) - - Function prediction TerpeneMiner + TTT 4e-4, 1e-3 2 2, 4, 8 30 - - Light attention + TTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
Table 4: Hyperparameters used for adapting TTT to individual models. The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction. Comma-separated lists show the values used for hyperparameter grid search, while the final values selected for computing the test results are highlighted in bold. Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone.
Appendix B Extended results
@@
In this section, we provide additional results on test sets (Section B.1) and discuss validation performance (Section B.2).
B.1 Detailed test performance
@@
In this section, we provide details on the test performance. Specifically, Table 5 shows that test-time training (TTT) primarily enhances performance on challenging targets, characterized by a low number of similar proteins in sequence databases, as measured by MSA depth. Additionally, we provide an example illustrating how TTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations within the protein core (Figure 8).
@@
Next, Figure 9 shows the distribution of TTT effects: in many cases, TTT has minimal impact on performance; often, it leads to substantial improvements; and in rare cases TTT results in a decrease in performance. This positions TTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research. While we demonstrate these effects using a protein folding example, we observe a similar distribution of TTT impact across the tasks.
@@
We also observe that the overall trend of TTT generally leads to improved performance, with robust consistency across random seeds. However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure 10).
B.2 Validation performance
@@
This section discusses the performance of test-time training (TTT) on validation data. Table 6 illustrates the validation performance of all tested methods for fitness prediction on our newly constructed MaveDB dataset. TTT enhances the performance of all the methods.
@@
The primary focus of the section is hyperparameter tuning. Table 4 provides the grid of hyperparameters explored for each model and its size. Figure 11 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single test protein. While most hyperparameter configurations lead to overall improvements when using TTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting. However, with a reliable predicted confidence measure, such as pLDDT, the appropriate TTT step can be selected to mitigate overfitting. Figure 12 demonstrates that when using ESM3 + TTT with pLDDT-based step selection for protein folding, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
Figure 8: Example of test-time training (TTT) applied to fitness prediction. Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left). ESM2 + TTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle). The ground-truth fitness data aligns with the TTT-enhanced model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right). Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.
[…]
[] Avg. Spearman $\uparrow$ Spearman by MSA depth $\uparrow$ Low depth Medium depth High depth ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451 ESM2 (35M) + TTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598 SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057 SaProt (35M) + TTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091 ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153 ESM2 (650M) + TTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075 SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448 SaProt (650M) + TTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
@@
Table 5: Test-time training (TTT) performance on ProteinGym depending on MSA depth. MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section A.1.2). TTT primarily improves performance on difficult targets, with low MSA depth. Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
[…]
[] Avg. Spearman $\uparrow$ ESM2 (35M) (Lin et al., 2023) 0.4458 ESM2 (35M) + TTT (Ours) 0.4593 ESM2 (650M) (Lin et al., 2023) 0.4568 ESM2 (650M) + TTT (Ours) 0.4604 SaProt (650M) (Su et al., 2023) 0.4926 SaProt (650M) + TTT (Ours) 0.4926 SaProt (35M) (Su et al., 2023) 0.5251 SaProt (35M) + TTT (Ours) 0.5271
@@
Table 6: Performance of test-time training (TTT) on the MaveDB dataset. In this work, we use our newly constructed MaveDB benchmark as a validation fold for tuning the hyper-parameters of TTT for fitness prediction. For computational efficiency, we only select a subset of 50 proteins (Section A.1.1) and do not run TTT across multiple random seeds to estimate standard deviations. The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
@@
Figure 9: Per-protein performance of ESMFold + TTT and ESM3 + TTT on the CAMEO test set. The y-axis shows the change in TM-score after applying test-time training (TTT), with higher values indicating improvement. The x-axis represents performance across five random seeds. The red dashed line marks no change in TM-score (TM-score difference $=0$), and the pink band represents minor changes in TM-score ($-0.05<$ TM-score difference $<0.05$), which we do not consider significant. Each point in the swarm plot corresponds to a single protein from the CAMEO test set. On average, applying TTT to ESMFold improves the structure predictions for 7 out of 18 proteins, with 2 showing degradation. The rest of the proteins are not significantly affected. Similarly, applying TTT to ESM3 results in 6 improvements out of 18 proteins, with 1 case of degradation.
@@
Figure 10: Test performance of TerpeneMiner + TTT across fine-tuning steps. The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
@@
Figure 11: Dependence on hyperparameters in test-time training for fitness prediction. Each plot shows the progression of Spearman correlation (green) increasing alongside a decrease in perplexity (pink) for each TTT step, averaged across all assays in the MaveDB validation dataset. The model used is ESM2 (35M) + TTT and the grid displays the progression for all hyperparameter combinations. As the learning rate increases (left to right) and the effective batch size grows, given by the increased number of gradient accumulation steps (top to bottom), the model reaches peak performance more quickly and begins overfitting to a test protein. The optimal hyperparameter combination (learning rate = 4e-4, gradient accumulation steps = 16) lies near the center of the grid, balancing between underfitting and overfitting to a test protein.
@@
Figure 12: Hyperparameter search for protein structure prediction with ESM3 + TTT. We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”). We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the beta30 and betalinear30 distributions proposed in the ESM3 paper (Hayes et al., 2024). Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold. The last row, denoted as “No TTT”, shows the performance of ESM3 without TTT. The results indicate that ESM3 + TTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations. We selected the configuration from the last row to compute the results on the test fold. For the hyperparameter search, we used 30 TTT steps instead of 50 to reduce computation time.

### v2

@@ E.3.3 Models
Light attention + ProteinTTT.
When applying ProteinTTT to Light attention, we treat the frozen ESM-1b as the backbone $f$, the language modeling head of ESM-1b as the self-supervised head $g$, and the Light attention block as the fine-tuning head $h$.
Table A3: Hyperparameters used for adapting ProteinTTT to individual models. The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction. Comma-separated lists show the values used for hyperparameter grid search, while the final values selected for computing the test results are highlighted in bold. Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone. Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A10. Therefore, we used 30 steps in all our experiments. The only exception was ESM3 + ProteinTTT, where the number of steps was set to 50 during initial experiments with different models/tasks conducted in parallel before standardizing the number of steps to 30. Methods marked with an asterisk (“*”) used a slightly different calculation of the loss function. Specifically, the loss was propagated over all tokens, including special and non-masking tokens, while averaging the loss across all tokens simultaneously, rather than first averaging over sequences. This approach was used in the early stages of development, and we provide it in our codebase via loss_kind = ‘‘unnormalized_cross_entropy’’. Please note that MSA Transformer always uses 1 MSA in a batch and the “Batch size” represents the number of sequences in this MSA with the target sequence always present as the first one.
Learning rate Batch size Grad. acc. steps Steps (Conf. func. $c$) LoRA rank $r$ LoRa $\alpha$
Fitness prediction
ESM2 (35M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32, 64 30 - -
ESM2 (650M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
SaProt (35M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
SaProt (650M) + ProteinTTT * 4e-5, 4e-4, 4e-3 2, 4 4, 8, 16, 32 30 - -
ProSST (K=2048) + ProteinTTT * 1e-5, 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
@@
MSA Transformer + ProteinTTT 1e-6, 3e-6, 1e-5, 3e-5, 1e-4 1, 4, 8, 16, 32 1, 2, 4, 8 30 - -
Structure prediction
ESMFold + ProteinTTT 4e-4 4 4, 8, 32, 64 30 (pLDDT) 4, 8, 32 8, 16, 32
HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
Function prediction
EnzymeExplorer + ProteinTTT 4e-4, 1e-3 2 2, 4, 8 30 - -
Light attention + ProteinTTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
Appendix F Case study details
F.1 Modeling antibody-antigen loops
We download the SAbDab dataset from the official website [^12] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(Dunbar et al., 2014). We apply ProteinTTT to targets with low-confidence ESMFold predictions (pLDDT < 70) and remove sequences longer than 400 residues due to GPU memory limitations. This results in a final set of 175 antibody and 814 antigen chains. We predict the full structures using ESMFold+ProteinTTT (with the same hyperparameters tuned on the CAMEO validation set specified in Table˜A3) and compute LDDT scores against the corresponding PDB structures to assess local errors, which are particularly relevant for loop regions. For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (Chothia & Lesk, 1987) and calculate LDDT on these regions.
F.2 Expanding known structures of viral proteins
We use BFVD version archived/2023_02_v2131313https://bfvd.steineggerlab.workers.dev. This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (Mirdita et al., 2022) with MSAs constructed using Logan (Chikhi et al., 2024), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan. In Figure˜5, we also report pLDDT values for BFVD version archived/2023_02_v1, where structures are simply obtained from ColabFold with MSAs from Logan, i.e., strategy (i). We re-predict structures using ESMFold and ESMFold+ProteinTTT for sequences with length < 450 due to GPU memory constraints. We use the same hyperparameters tuned on the CAMEO validation set, as specified in Table˜A3, with the exception of 20 instead of 30 steps for computational efficiency.
Appendix G Extended results
@@
In this section, we provide additional results on test sets (Section˜G.1), discuss validation performance (Section˜G.2), and analyze the runtime performance of customization (Section˜G.3).
G.1 Detailed test performance
@@
In this section, we provide details on the test performance. Specifically, Table˜A4 shows that customization with ProteinTTT primarily enhances performance on challenging targets, characterized by a low number of similar proteins in sequence databases, as measured by MSA depth. Additionally, we provide a qualitative example illustrating how ProteinTTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations in the protein core (Figure A5).
@@
Next, Figure˜A6 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance. This positions ProteinTTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research. While we demonstrate these effects using a protein folding example, we observe a similar distribution of ProteinTTT impact across the tasks.
@@
We also observe that the overall trend of customization with ProteinTTT generally leads to improved performance, with robust consistency across random seeds. However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A8).
G.2 Validation performance
@@
This section discusses the performance of ProteinTTT on validation data. Table˜A5 illustrates the validation performance of the tested methods for fitness prediction on our newly constructed MaveDB dataset. ProteinTTT enhances the performance of all the methods.
@@
Next, we discuss the hyperparameter optimization. Table˜A3 provides the grid of hyperparameters explored for each model and its size, as well as specifies the optimal hyperparameters suitable for downstream applications. Figure˜A10 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein. While most of reasonable hyperparameter configurations lead to overall improvements when using customization with ProteinTTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting. However, with a reliable predicted confidence measure, such as pLDDT, the appropriate customization step for each protein can be selected to mitigate overfitting. Figure˜A11 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
G.3 Runtime performance
In this section, we demonstrate that customization with ProteinTTT can be done efficiently, with an acceptable computational overhead. Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (Jumper et al., 2021) or AlphaFold3 (Abramson et al., 2024), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
This observation highlights the practical utility of ProteinTTT. For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (Lin et al., 2023). Nevertheless, the original ESMFold has high confidence predictions only for 36% of sequences from the metagenomic database, while the other 392 million sequences remain with low or medium confidence predictions. At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A6 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences). It means that applying ESMFold + ProteinTTT to these remaining sequences could significantly expand the metagenomic atlas characterized by ESMFold. Here, we illustrate this on a similar case study by applying ESMFold + ProteinTTT to more than 300 thousand viral proteins in BFVD (Section˜5.2)
@@ G.3 Runtime performance
Table A4: ProteinTTT performance on ProteinGym depending on MSA depth. MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜E.2.2). Customization with ProteinTTT primarily improves performance on difficult targets, with low MSA depth. Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
Avg. Spearman $\uparrow$
Spearman by MSA depth $\uparrow$
Low depth Medium depth High depth
ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451
ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598
SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057
SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091
ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153
ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075
SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448
SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
ProSST (K=2048) (Li et al., 2024) 0.5068 0.4731 0.5107 0.5749
ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
@@
Table A5: Performance of ProteinTTT on the MaveDB dataset. In this work, we use our newly constructed MaveDB dataset as a validation fold for tuning the ProteinTTT hyper-parameters for fitness prediction. For computational efficiency, we only select a subset of 50 proteins (Section˜E.2.1) and do not run customization across multiple random seeds to estimate standard deviations. The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
Avg. Spearman $\uparrow$
ESM2 (35M) (Lin et al., 2023) 0.4458
ESM2 (35M) + ProteinTTT (Ours) 0.4593
ESM2 (650M) (Lin et al., 2023) 0.4568
ESM2 (650M) + ProteinTTT (Ours) 0.4604
SaProt (650M) (Su et al., 2023) 0.4926
SaProt (650M) + ProteinTTT (Ours) 0.4926
SaProt (35M) (Su et al., 2023) 0.5251
SaProt (35M) + ProteinTTT (Ours) 0.5271
ProSST (K=2048) (Li et al., 2024) 0.5444
ProSST (K=2048) + ProteinTTT (Ours) 0.5462
Figure A4: Running time of ESMFold + ProteinTTT. For ESMFold and its variants, the median and interquartile ranges of running times on the CAMEO test set are shown using a single NVIDIA A100 GPU. For AlphaFold2, we use estimates from Lin et al. (2023). Specifically, a forward pass through AlphaFold2 is approximately 60 times more computationally expensive than ESMFold (e.g., AlphaFold2 (no MSA; estimate): $2\times 60=120$ seconds), with additional MSA construction taking at least 10 minutes using standard pipelines (AlphaFold2 (estimate): $2\times 60+10\times 60=720$ seconds). ESMFold + ProteinTTT (30 steps) involves LoRA parameter updates, along with forward passes at each customization step to estimate pLDDT and select the structure with the highest predicted confidence. Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜G.2). Overall, ESMFold + ProteinTTT maintains the speed advantage of ESMFold, and is at least an order of magnitude faster than AlphaFold2.
Figure A5: Example of protein fitness prediction upon single-sequence model customization with ProteinTTT. Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left). ESM2 + ProteinTTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle). The ground-truth fitness data aligns with the customized model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right). Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.
@@
Figure A6: Per-protein performance of ESMFold + ProteinTTT and ESM3 + ProteinTTT on the CAMEO test set. The y-axis shows the change in TM-score after applying customization with ProteinTTT, with higher values indicating improvement. The x-axis represents performance across five random seeds. The red dashed line marks no change in TM-score (TM-score difference =0=0), and the pink band represents minor changes in TM-score (0.05<-0.05< TM-score difference <0.05<0.05), which we do not consider significant. Each point in the swarm plot corresponds to a single protein from the CAMEO test set. On average, applying ProteinTTT to ESMFold improves the structure predictions for 7 out of 18 proteins, with 2 showing degradation. The rest of the proteins are not significantly affected. Similarly, applying ProteinTTT to ESM3 results in 6 improvements out of 18 proteins, with 1 case of degradation.
Figure A7: Test performance of ESMFold + ProteinTTT and ESM3 + ProteinTTT on the CAMEO test set depending on the total number of customization steps. The x-axis shows the averaged performance across all test proteins, with error bars representing the standard deviation across five random seeds. The y-axis metrics correspond to the structure with the highest pLDDT score up to the given step. While an increased number of ProteinTTT steps generally enhances performance, only a few steps (e.g., five) may suffice to achieve significant performance improvement.
@@
Figure A8: Test performance of EnzymeExplorer + ProteinTTT across customization steps. The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
Figure A9: ESMFold + ProteinTTT pLDDT correlates with ESMFold + ProteinTTT LDDT. The evaluation was performed on 17,582 AlphaFold2 reference structures from the BFVD database with pLDDT > 90. Here, $r=0.875$ denotes the Pearson correlation coefficient.
@@
Figure A10: Dependence on ProteinTTT hyperparameters for customized fitness prediction. Each plot shows the progression of Spearman correlation (green) increasing alongside a decrease in perplexity (pink) for each customization step, averaged across all assays in the MaveDB validation dataset. The model used is ESM2 (35M) + ProteinTTT, and the grid displays the combinations of different numbers of gradient accumulation steps (i.e., effective batch sizes; shown in rows, increasing from top to bottom) and learning rates (columns, increasing from left to right). As the learning rate increases and the number of gradient accumulation steps grows, the model reaches peak performance more quickly but begins to overfit to a target protein. The optimal hyperparameter combination (learning rate = 4e-4, gradient accumulation steps = 16) lies near the center of the grid, balancing between underfitting and overfitting to a target protein. Notably, the figure demonstrates that, although ProteinTTT involves three main hyperparameters (batch size, learning rate, and the number of steps), there are effectively only two degrees of freedom controlling the performance of the model. In other words, by keeping the number of steps constant (e.g., 30), the expected performance can be controlled by adjusting the learning rate and the batch size.
@@
Figure A11: Hyperparameter search for protein structure prediction with ESM3 + ProteinTTT. We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”). We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (Hayes et al., 2024). Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold. The last row, denoted as “No ProteinTTT”, shows the performance of ESM3 without customization. The results indicate that ESM3 + ProteinTTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations. We selected the configuration from the last row (excluding “No ProteinTTT”) to compute the results on the test fold. For the hyperparameter search, we used 30 customization steps instead of 50 to reduce computation time.